# The Central Role of the Loss Function in Reinforcement Learning

### v1

@@ 1 Introduction
The value-based approach to reinforcement learning (RL) reduces the decision making problem to regression: first predict the expected rewards to go under the optimal policy, given state and action, and then one can simply choose the action that maximizes the prediction at every state. This regression, called $Q$-learning [50], combined with recent decades’ advances in deep learning, plays a central role in the empirical successes of deep RL. A prime example is DeepMind’s groundbreaking use in 2014 of deep $Q$-networks to play Atari with no feature engineering [36].
In prediction, we often say a good model is one with low mean-squared error out of sample. Correspondingly, regression is usually done by minimizing the average squared loss between predictions and targets in the training data. However, low mean-squared error may translate only loosely to good downstream decision making. Is squared loss the right choice for learning $Q$-functions?
In this article, we highlight that the answer to this question is a resounding “no.” Across both offline and online RL, alternative loss functions work better both empirically and theoretically. In this article we focus on the theoretical question: when and how do alternative loss functions attain better guarantees for decision making? A recent flurry of papers give regret upper bounds that adapt to special (although practically common) settings, such as low optimal expected costs or low returns variance, where they attain faster rates. We explain the phenomenon in a hopefully elucidating manner, starting with the simple setting of cost-sensitive classification and then building up to reinforcement learning. The technical material is largely based on Wang et al. [46, 47], Ayoub et al. [4], Foster and Krishnamurthy [18], with a couple new results along the way.
@@ 2.1 Solving CSC with Squared-Loss Regression
A value-based approach to CSC is to learn a cost prediction $f(x,a)\approx\bar{C}(x,a)$ by regressing costs on contexts and then use an induced policy: $\pi_{f}(x)\in\mathop{\rm arg\,min}_{a}f(x,a)$. A standard way to learn such a cost prediction is to minimize squared error. To see why this yields a good policy, define the squared loss as
$\ell_{\operatorname{sq}}(\hat{y},y):=(\hat{y}-y)^{2}$
Define the excess squared-loss risk of a prediction $f$ as $\mathcal{E}_{\operatorname{sq}}(f):=\sum_{a}\mathbb{E}[\ell_{\operatorname{sq}}(f(x,a),c(a))-\ell_{\operatorname{sq}}(\bar{C}(x,a),c(a))]$. This straightforwardly bounds the suboptimality of its induced policy:
@@ 2.2 The Second-Order Lemma
The PAC guarantee above shrinks at a nice parametric rate of $\mathcal{O}(n^{-1/2})$ as the number of samples $n$ grows, but can we do better? The bound in Eq. 1, which translates error in predicted means to excess risk in our loss function, was rather loose.
@@
Here, $h^{2}(p,q)$ is the squared Hellinger distance, which is an $f$-divergence, and it is bounded in $[0,1]$. This lemma is equivalent to Lemma 4.3 of [47] and we provide a simplified proof in Sec. 2.6. Interpreting the inequality, which is a completely deterministic statement, in terms of estimating means, it says that estimation error can be bounded by two terms: one involves the standard deviation times a discrepancy and the other is a squared discrepancy. As variance shrinks, the first term vanishes and the second term dominates, which as a squared term we expect to decay quickly.
@@ 2.3 Regression with the Binary-Cross-Entropy Loss: First-Order PAC Bounds for CSC
where $h^{2}_{\operatorname{Ber}}(f,g)=\frac{1}{2}(\sqrt{f}-\sqrt{g})^{2}+\frac{1}{2}(\sqrt{1-f}-\sqrt{1-g})^{2}$ is the squared Hellinger distance between Bernoullis with means $f$ and $g$. This recovers the key inequalities in Wang et al. [46], Foster and Krishnamurthy [18], Ayoub et al. [4].
@@
This inequality also means that we could use the excess bce-loss risk to bound Eq. 1. The point of using bce loss, however, is to do better than Eq. 1 via Eq. 2.
@@
Thus, together with Eq. 3 and Lem. 2, we have shown the following PAC bound for bce-loss regression:
@@
This result was first proved in Theorem 3 of [18]. Notably, the bound is adaptive to the optimal expected costs $V^{\star}$ and converges at a fast $n^{-1}$ rate when $V^{\star}\lesssim 1/n$. Under the cost minimization setup, first-order bounds are also called ‘small-cost’ bounds since they converge at a fast rate when the optimal cost $V^{\star}$ is small.
@@ 2.4 Maximum Likelihood Estimation: Second-Order PAC Bounds for CSC
Since costs are bounded in $[0,1]$, we observe that $\sigma^{2}(\pi)\leq V(\pi)$, and hence a second-order bound is tighter than a first-order bound. We note however that our approach is distributional and the bounds depend on $\ln|\mathcal{P}|$ which might be larger than $\ln|\mathcal{F}|$; also, Assump. 2 may be more stringent than Assump. 1, although in practice, distributional approaches still often achieve superior performance [46, 47, 7].
@@ 2.5 Improved Second-Order PAC Bounds for CSC with Pessimistic MLE
Proof.
For both claims, we condition on Eq. 5 which holds w.p.a.l. $1-\delta$. For Claim (a): for any $p\in\mathcal{P}_{n}$ (which includes $\hat{p}^{\operatorname{pes}}$), we have $n\mathcal{E}_{\operatorname{mle}}(p)\leq\frac{1}{2}L_{\operatorname{mle}}(p)-\frac{1}{2}L_{\operatorname{mle}}(\hat{p}^{\operatorname{mle}}_{\mathcal{P}})+A\ln(A|\mathcal{P}|/\delta)\leq\frac{1}{2}\beta+A\ln(A|\mathcal{P}|/\delta)\leq 2A\ln(A|\mathcal{P}|/\delta)$, where the first inequality is by Eq. 5 and the fact that $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$ minimizes the empirical risk; and the second inequality is by the definition of $\mathcal{P}_{n}$. To prove Claim (b), we first show that $C\in\mathcal{P}_{n}$: by Eq. 5 and the non-negativity of $\mathcal{E}_{\operatorname{mle}}$, we have $L_{\operatorname{mle}}(C)-L_{\operatorname{mle}}(p)\leq 2A\ln(A|\mathcal{P}|/\delta)=\beta$ for all $p\in\mathcal{P}$ (which includes $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$). Thus, this shows that $C$ satisfies the $\mathcal{P}_{n}$ condition, implying its membership in the set. To conclude Claim (b), we have $\sum_{i=1}^{n}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},\pi^{\star}(x_{i}))\geq\sum_{i=1}^{n}\min_{a}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},a)\geq\sum_{i=1}^{n}\min_{a}\bar{C}(x_{i},a)$. Claim (b) then follows by multiplicative Chernoff [54, Theorem 13.5]. ∎
@@
Finally, we note that pessimistic mle requires more computation than plug-in mle, since we have the extra step of optimizing over $\mathcal{P}_{n}$. For one-step settings like CSC or contextual bandits, this can be feasibly implemented with binary search [19] or width computation [16]. However, in multi-step settings like RL as we will soon see, this optimization problem is NP-hard [13].
@@ 2.6 Proof of the Second-Order Lemma
which is a simple consequence of Cauchy-Schwartz [46, Lemma A.1]. Thus, we first prove the second-order lemma using ($\triangle$($\cdot$)△ ( ⋅ ), which is more natural, and then convert the bounds to $h^{2}(\cdot)$ using Eq. 9.
@@ 3.1 Plug-in Squared Loss Cannot Achieve First-Order
Probability of the bad event. For $\mathfrak{E}_{1}$, since $n(x^{2})\sim\operatorname{Bin}(n,n^{-1})$, thus $\Pr(n(x^{2})=1)=(1-n^{-1})^{n-1}\geq e^{-1}$. Hence, $\Pr(\mathfrak{E}_{1})\geq(2e)^{-1}$. For $\mathfrak{E}_{2}$, we apply the multiplicative Chernoff bound [54, Theorem 13.5], which implies $\widehat{\mu}(x^{1},a^{1})<2\mu_{n}+\frac{3}{n-1}$ w.p.a.l. $1-e^{-3}$. Thus, $\Pr(\mathfrak{E}_{1}\cap\mathfrak{E}_{2})\geq 1-(1-(2e)^{-1})-e^{-3}\geq 0.1$. ∎
@@ 3.2 Plug-in BCE Loss Cannot Achieve Second-Order
Next, we show that the bce-loss induced policy cannot achieve second-order bounds. This is a new result.
@@ 4.1 Problem Setup
Bellman Equations. We recall the Bellman equations. Let $\mathcal{T}^{\pi}$ denote the Bellman operator for policy $\pi$, defined by $\mathcal{T}_{h}^{\pi}f(x,a)=\mathbb{E}_{c\sim C_{h}(x,a),x^{\prime}\sim P_{h}(x,a)}[c+f(x^{\prime},\pi_{h+1})]$ for any function $f$. The Bellman equations are $f_{h}=\mathcal{T}^{\pi}_{h}f_{h+1}$ for all $h$, where $f_{h}=Q^{\pi}_{h}$ is the unique solution. Also, the Bellman optimality operator $\mathcal{T}^{\star}$ is defined by $\mathcal{T}_{h}^{\star}f(x,a)=\mathbb{E}_{c\sim C_{h}(x,a),x^{\prime}\sim P_{h}(x,a)}[c+\min_{a^{\prime}}f(x^{\prime},a^{\prime})]$. The Bellman optimality equations are $f_{h}=\mathcal{T}^{\star}_{h}f_{h+1}$ for all $h$, where $f_{h}=Q^{\star}_{h}$ is the unique solution.
@@ 4.2 Solving Online RL with Optimistic Squared-Loss Regression
This shows that Alg. 2 with the squared loss is guaranteed to learn a policy that converges to the optimal policy at a $\Theta(K^{-1/2})$ rate, which is the minimax-optimal rate. In Sec. 4.3 we show that the V-type dimension can be bounded by the rank of the transition kernel in a low-rank MDP [2], a canonical model for RL with non-linear function approximation.
Computation of Version Space Algorithms. While we presented version space algorithms for their simplicity and statistical efficiency, we note here that they are computationally hard to run in general. Specifically, the computational cost of optimizing over the version space is NP-hard even in tabular MDPs [13]. However, in the one-step $H=1$ setting (a.k.a. contextual bandits), the version space optimization is oracle-efficient [19, 16, 47]. In the RL setting, there are also approaches to mitigate the computational hardness of optimism. One approach is to use $\varepsilon$-greedy as a computationally efficient but more myopic exploration strategy – this has been successful in practice [36, 6] and also enjoys theoretical guarantees under assumptions about the easiness of exploration [14, 55]. Another approach is to assume access to an offline dataset that already has good coverage and so strategic exploration is no longer necessary. This setting is called hybrid RL [41] and we revisit this in Sec. 4.7.
We now prove Thm. 7.
@@ 4.5 Second-Order Bounds for Online RL with Optimistic MLE: Benefits of Distributional RL
The above mle loss bounds scales with the variances of the policies selected by the algorithm, and are thus are called second-order (a.k.a. variance dependent) bounds. As we saw in the CSC setting, a second-order bound is actually strictly sharper than the first-order bound and this is also true in RL [47, Theorem 2.1]. The variance bound can be much tighter in near-deterministic settings where the optimal policy’s cost is far from zero.
While DistRL achieves the tightest scaling w.r.t. $K$ in this paper, it does have a drawback compared to the bce result Thm. 8: DistRL requires modeling full distributions while the bce loss only requires modeling the mean. Specifically, DistBC is a stronger requirement than the standard BC, and the conditional distribution class $\mathcal{P}$ is in general larger than the regression class $\mathcal{F}$. Nevertheless, in low-rank MDPs with discrete cost distributions, [47, Section 5.1] proved that a linear conditional distribution class both satisfies DistBC and has a bracketing entropy of $\widetilde{\mathcal{O}}(dM+\log|\Phi|)$ where $M$ is the number of discretizations. Also, deep RL algorithms such as C51 [6] and IQN [12] have been shown to effectively learn the cost-to-go distributions, which suggests that distributional TD learning is feasible with expressive neural networks [7].
Proof.

### v2

@@ 1 Introduction
The value-based approach to reinforcement learning (RL) reduces the decision making problem to regression: first predict the expected rewards to go under the optimal policy, given state and action, and then one can simply choose the action that maximizes the prediction at every state. This regression, called $Q$-learning [67], combined with recent decades’ advances in deep learning, plays a central role in the empirical successes of deep RL. Groundbreaking examples are DeepMind’s use of deep $Q$-networks to play Atari with no feature engineering [48] and OpenAI’s use of deep reward models to align large language models with human preferences via RL fine-tuning [51].
In prediction, we often say a good model is one with low mean-squared error out of sample. Correspondingly, regression is usually done by minimizing the average squared loss between predictions and targets in the training data. However, low mean-squared error may translate loosely to high-quality, downstream decision making. Thus, the natural question arises: is squared loss always the best choice for learning $Q$-functions?
@@
In this survey article, we highlight that the answer to this question is a resounding “no!” We focus on the theoretical question: when and how do alternative loss functions attain better guarantees for decision making? We start with the simple setting of cost-sensitive classification (CSC) and show that the binary cross-entropy (bce) loss leads to an improved convergence guarantee that adapts to problem instances with low optimal cost; such a result is called a first-order bound. This result is due to Foster and Krishnamurthy [22], who first observed that the bce loss can have benefits over square loss when outcomes are heteroskedastic and used this observation to derive first-order regret bounds for CSC and contextual bandits (CB). Then, we prove that the maximum likelihood (mle) loss yields an even better convergence guarantee that additionally adapts to problem instances with low variance; such a result is called a second-order bound and is strictly stronger than first-order. This result is due to Wang et al. [62], who first observed that the mle loss can have benefits over the bce loss when outcomes have low variance and used this observation to derive second-order bounds for CB and RL. We also provide lower bounds to show that the separation between these loss functions is indeed real, with one result new to this paper separating bce and mle in their ability to attain second-order bounds. The decision making performance that each loss function attains in each type of problem instance is outlined in Table 1.
We then turn to extending these observations and intuitions to RL, following the results of Wang et al. [61, 62], Ayoub et al. [7]. We systematically derive bounds for both online and offline RL, revealing that the trends in Table 1 continue to hold for the more challenging RL settings. Finally, we discuss issues surrounding computational complexity and provide a solution in the hybrid RL setting [55]. In summary, we provide a detailed survey on the decision performance of different loss functions, with the aim of elucidating their central role in RL and data-driven decision making generally. The technical material is largely based on Foster and Krishnamurthy [22], Wang et al. [61, 62], Ayoub et al. [7] with a couple new results along the way.
@@ 2.1 Solving CSC with Squared-Loss Regression
A value-based approach to CSC is to learn a cost prediction $f(x,a)\approx\bar{C}(x,a)$ by regressing costs on contexts and then use an induced policy: $\pi_{f}(x)\in\mathop{\rm arg\,min}_{a}f(x,a)$. A standard way to learn such a cost prediction is to minimize squared error [5, 43, 28].
Define the squared loss as
$\ell_{\operatorname{sq}}(\hat{y},y):=(\hat{y}-y)^{2},$
and define the excess squared-loss risk of a prediction $f$ as $\mathcal{E}_{\operatorname{sq}}(f):=\sum_{a}\mathbb{E}[\ell_{\operatorname{sq}}(f(x,a),c(a))-\ell_{\operatorname{sq}}(\bar{C}(x,a),c(a))]$. This straightforwardly bounds the suboptimality of its induced policy:
@@
This PAC bound shrinks at a nice parametric rate of $\mathcal{O}(n^{-1/2})$ as the number of samples $n$ grows, but can we do better? The step in Eq. 1, which translates error in predicted means to excess risk in the squared loss, was rather loose. Foster and Krishnamurthy [22] thus investigated the bce loss and showed that it can achieve a first-order bound, which we present next after introducing an important second-order lemma.
@@ 2.2 The Second-Order Lemma
Here, $h^{2}(p,q)$ is the squared Hellinger distance, which is an $f$-divergence, and it is bounded in $[0,1]$. This lemma is equivalent to Lemma 4.3 of [62] and we provide a simplified proof in Sec. 2.6. Interpreting the inequality, which is a completely deterministic statement, in terms of estimating means, it says that estimation error can be bounded by two terms: one involves the standard deviation times a discrepancy and the other is a squared discrepancy. As variance shrinks, the first term vanishes and the second term dominates, which as a squared term we expect to decay quickly. We note [24, 25] are related and provide first-moment and second-moment bounds, which do not directly imply the tighter variance bound in Lem. 1.
@@ 2.3 Regression with the Binary-Cross-Entropy Loss: First-Order PAC Bounds for CSC
where $h^{2}_{\operatorname{Ber}}(f,g)=\frac{1}{2}(\sqrt{f}-\sqrt{g})^{2}+\frac{1}{2}(\sqrt{1-f}-\sqrt{1-g})^{2}$ is the squared Hellinger distance between Bernoullis with means $f$ and $g$. This recovers the key inequalities in [22, 61, 7].
@@
This inequality also means that we could use the excess bce-loss risk to bound Eq. 1 in place of excess squared-loss risk. The point of using bce loss, however, is to do better than Eq. 1 via Eq. 2.
@@
Thus, together with Eq. 3 and Lem. 2, we have shown a first-order PAC bound for bce-loss regression:
@@
This result was first observed in Theorem 3 of [22]. Notably, the bound is adaptive to the optimal expected costs $V^{\star}$ and converges at a fast $n^{-1}$ rate when $V^{\star}\lesssim 1/n$. Under the cost minimization setup, first-order bounds are also called ‘small-cost’ bounds since they converge at a fast rate when the optimal cost $V^{\star}$ is small.
@@ 2.4 Maximum Likelihood Estimation: Second-Order PAC Bounds for CSC
Since costs are bounded in $[0,1]$, we observe that $\sigma^{2}(\pi)\leq V(\pi)$, and hence a second-order bound is tighter than a first-order bound. However, our approach is distributional with bounds depending on $\ln|\mathcal{P}|$ which can be larger than $\ln|\mathcal{F}|$; also, Assump. 2 may be more stringent than Assump. 1, although in practice, distributional approaches still achieve superior performance [61, 62, 10].
@@ 2.5 Improved Second-Order PAC Bounds for CSC with Pessimistic MLE
Proof of Lem. 5.
For both claims, we condition on Eq. 5 which holds w.p.a.l. $1-\delta$. For Claim (a): for any $p\in\mathcal{P}_{n}$ (which includes $\hat{p}^{\operatorname{pes}}$), we have
$\textstyle n\mathcal{E}_{\operatorname{mle}}(p)\leq\frac{1}{2}L_{\operatorname{mle}}(p)-\frac{1}{2}L_{\operatorname{mle}}(\hat{p}^{\operatorname{mle}}_{\mathcal{P}})+A\ln(A|\mathcal{P}|/\delta)$
$\textstyle\leq\frac{1}{2}\beta+A\ln(A|\mathcal{P}|/\delta)\leq 2A\ln(A|\mathcal{P}|/\delta),$
where the first inequality is by Eq. 5 and the fact that $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$ minimizes the empirical risk; and the second inequality is by the definition of $\mathcal{P}_{n}$. To prove Claim (b), we first show that $C\in\mathcal{P}_{n}$: by Eq. 5 and the non-negativity of $\mathcal{E}_{\operatorname{mle}}$, we have $L_{\operatorname{mle}}(C)-L_{\operatorname{mle}}(p)\leq 2A\ln(A|\mathcal{P}|/\delta)=\beta$ for all $p\in\mathcal{P}$ (which includes $\hat{p}^{\operatorname{mle}}_{\mathcal{P}}$). Thus, this shows that $C$ satisfies the $\mathcal{P}_{n}$ condition, implying its membership in the set. To conclude Claim (b), we have $\sum_{i=1}^{n}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},\pi^{\star}(x_{i}))\geq\sum_{i=1}^{n}\min_{a}\overline{\hat{p}^{\operatorname{pes}}}(x_{i},a)\geq\sum_{i=1}^{n}\min_{a}\bar{C}(x_{i},a)$. Claim (b) then follows by multiplicative Chernoff [71, Theorem 13.5]. ∎
@@
Finally, we note that pessimistic mle requires more computation than plug-in mle, since we have the extra step of optimizing over $\mathcal{P}_{n}$. For one-step settings like CSC or contextual bandits, this can be feasibly implemented with binary search [23] or width computation [39, 20]. However, in multi-step settings like RL as we will soon see, this optimization problem is NP-hard [17].
@@ 2.6 Proof of the Second-Order Lemma
which is a simple consequence of Cauchy-Schwartz [61, Lemma A.1]. Thus, we first prove the second-order lemma using ($\triangle$($\cdot$)△ ( ⋅ ), which is more natural, and then convert the bounds to $h^{2}(\cdot)$ using Eq. 9.
@@ 3.1 Plug-in Squared Loss Cannot Achieve First-Order
Probability of the bad event. For $\mathfrak{E}_{1}$, since $n(x^{2})\sim\operatorname{Bin}(n,n^{-1})$, thus $\Pr(n(x^{2})=1)=(1-n^{-1})^{n-1}\geq e^{-1}$. Hence, $\Pr(\mathfrak{E}_{1})\geq(2e)^{-1}$. For $\mathfrak{E}_{2}$, we apply the multiplicative Chernoff bound [71, Theorem 13.5], which implies $\widehat{\mu}(x^{1},a^{1})<2\mu_{n}+\frac{3}{n-1}$ w.p.a.l. $1-e^{-3}$. Thus, $\Pr(\mathfrak{E}_{1}\cap\mathfrak{E}_{2})\geq 1-(1-(2e)^{-1})-e^{-3}\geq 0.1$. ∎
@@ 3.2 Plug-in BCE Loss Cannot Achieve Second-Order
Next, we show that the bce-loss induced policy cannot achieve second-order bounds. This is a new result that did not appear before.
@@ 4.1 Problem Setup
Bellman equations. Define the Bellman operator $\mathcal{T}^{\pi}$ by $\mathcal{T}_{h}^{\pi}f(x,a)=\mathbb{E}_{c\sim C_{h}(x,a),x^{\prime}\sim P_{h}(x,a)}[c+f(x^{\prime},\pi_{h+1})]$ for any function $f$ and policy $\pi$ The Bellman equations are $f_{h}=\mathcal{T}^{\pi}_{h}f_{h+1}$ for all $h$, where $f_{h}=Q^{\pi}_{h}$ is the unique solution. Also, the Bellman optimality operator $\mathcal{T}^{\star}$ is defined by $\mathcal{T}_{h}^{\star}f(x,a)=\mathbb{E}_{c\sim C_{h}(x,a),x^{\prime}\sim P_{h}(x,a)}[c+\min_{a^{\prime}}f(x^{\prime},a^{\prime})]$. The Bellman optimality equations are $f_{h}=\mathcal{T}^{\star}_{h}f_{h+1}$ for all $h$, where $f_{h}=Q^{\star}_{h}$ is the unique solution.
@@ 4.2 Solving Online RL with Optimistic Squared-Loss Regression
This shows that Alg. 2 with the squared loss is guaranteed to learn a policy that converges to the optimal policy at a $\Theta(K^{-1/2})$ rate, which is the minimax-optimal rate. In Sec. 4.3 we show that the V-type dimension can be bounded by the rank of the transition kernel in a low-rank MDP [3, 4], a canonical model for RL with non-linear function approximation.
Computational complexity of version space algorithms. While we presented version space algorithms for their simplicity and statistical efficiency, we note here that they are computationally hard to run in general. Specifically, the computational cost of optimizing over the version space is NP-hard even in tabular MDPs [17]. However, in the one-step $H=1$ setting (a.k.a. contextual bandits), the version space optimization is oracle-efficient [23, 20, 62]. In the RL setting, there are also approaches to mitigate the computational hardness of optimism. One approach is to use $\varepsilon$-greedy as a computationally efficient but more myopic exploration strategy – this has been successful in practice [48, 9] and also enjoys theoretical guarantees under assumptions about the easiness of exploration [18, 72]. Another approach is to assume access to an offline dataset that already has good coverage and so strategic exploration is no longer necessary. This setting is called hybrid RL [55] and we revisit this in Sec. 4.7.
@@ 4.5 Second-Order Bounds for Online RL with Optimistic MLE: Benefits of Distributional RL
The above mle loss bounds scales with the variances of the policies selected by the algorithm, and are thus are called second-order (a.k.a. variance dependent) bounds. As we saw in the CSC setting, a second-order bound is actually strictly sharper than the first-order bound and this is also true in RL [62, Theorem 2.1]. The variance bound can be much tighter in near-deterministic settings where the optimal policy’s cost is far from zero.
While DistRL achieves the tightest scaling w.r.t. $K$ in this paper, it does have one drawback compared to the bce result (Thm. 8): DistRL requires modeling full distributions while the bce loss only requires modeling the mean. Specifically, DistBC is a stronger requirement than the standard BC, and the conditional distribution class $\mathcal{P}$ is in general larger than the regression class $\mathcal{F}$. Nevertheless, in low-rank MDPs with discrete cost distributions, this assumption can always be satisfied with a linear function class of distributions [62, Section 5.1], which also has bounded bracketing entropy of $\widetilde{\mathcal{O}}(dM+\log|\Phi|)$ where $M$ is the number of discretizations. Moreover, distributional TD may not be so restrictive beyond low-rank MDPs, since deep RL approaches with non-linear neural networks have been quite effective for estimating the cost-to-go distributions [9, 16, 10]. Closing this gap between theory and practice is an interesting direction.
We also highlight two important recent works on second-order bounds for contextual bandits, the one-step special case of RL where DistBC simplifies to distributional realizability (Assump. 2). In this setting, Pacchiano [52] proved a second-order bound for contextual bandits with only mean realizability, which is weaker than distributional realizability, by using thresholded least squares. Concurrently, Jia et al. [31] obtained a similar bound as [52] for the strong adversary setting and a complementary bound for the weak adversary setting. Jia et al. [31] also furnished a lower bound proving that the eluder-based second-order bounds in Thm. 9 and [62] are tight and unimprovable even if the number of actions is less than the eluder dimension.
Proof of Thm. 9.